A bank has ​$16 billion in total​ assets, of which ​$150 million are its total reserves. Customers initially hold a total of ​$9 billion in their demand deposit accounts with this​ bank, but the bank automatically transforms ​$8 billion of this amount into money market deposit accounts for those holding a large amount in their accounts. The desired reserve ratio for demand deposits is

15percent. The bank has no other reserve requirement for money market accounts.

Calculate this​ bank's desired reserve​ requirement: ​$____million. ​(Round your response to the nearest​ integer.)

Calculate the amount of excess reserves for this​ bank: ​$ ___million. ​(Round your response to the nearest​ integer.)

Calculate the amount of this​ bank's contribution to the M1 measure of​ money: ​$___billion. ​(Round your response to the nearest​ integer.)
Answer #1

Demand Deposit after $8 billion is transformed into money market deposit accounts = 9-8=$1 billion

Reserves Requirement = 15%*Demand Deposits =15%*1 billion = $150 million

Excess Reserves = Actual Reserves-Required Reserves = 150-150=$0 million

M1 only includes Demand Deposits hence contribution to M1 = $150 million